Find the integral surface of the partial differential equation
(x - y) y^2 p + (y - x) x^2 q = (x^2 + y^2) z
Through the curve xz = a^2 , y = 0
1
Expert's answer
2020-03-13T10:23:32-0400

dx/((x-y)*y2)=dy/((y-x)*x2)=dz/((x2+y2)*z)

"\int" y2dy = "\int"-x2dx

y3=-x3+c1 (c1 - constant) (1)

"\int" y2/z*dz = "\int" (x2+y2)/(x-y)dx

y2ln(z)=1/2*(x2-3y2+2xy+4y2ln(x-y))+c2 (c2 - constant) (2)

y3+x3 = c1

y2ln(z) - 1/2*(x2-3y2+2xy+4y2ln(x-y)) = c2


y=0

z = a2/x


c1 = x3

x = c11/3

c2 = - 1/2*(x2)

x = (-2*c2)1/2

c11/3 = (-2*c2)1/2

c12 = (-2*c2)3


x6=(-2*(y2ln(z) - 1/2*(x2-3y2+2xy+4y2ln(x-y))))3